40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3314     Entry time: Wed Jul 28 18:24:57 2010
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: 2 Green Periscopes have mirrors, aligned 

[Koji, Jenne, Kiwamu]

This is to describe the work that went on in the Cleanroom today.  Kiwamu's entry will detail the tidbits that happened in the chamber.

We engraved the periscope mounts with the mirror info for the mirrors which were placed in the periscope.  We also engraved the barrels of the optics with their info, for posterity.  Koji carefully put the mirrors into the periscopes.  Since we have wedged optics, the goal was to have the front HR surface of the mirror parallel to the plane of the mount, and leave a bit of space behind one side of the optic (if we just pushed the optic fully in, the HR surface wouldn't be flat, and would send the beam off to the left or right somewhere).  Once the mirrors were mounted in the periscopes, we checked the vertical levelness of the outcoming beam.  For the first periscope (the one which has been installed on the BS table), the beam was deflected upward (2.5)/32 inches over 55inches.  This corresponds to a 1.4mRad vertical deflection.  The second periscope (which will eventually be installed on the OMC table) had a deflection of 1/32 over 55inches, or 0.6mRad.  We did not check the side-to-side deflection for either of the periscopes.

We also engraved one more DLC mount with mirror info, and put a mirror into the mount.  This is one of the optics that was placed onto the BS table today, which Kiwamu will describe.

We removed TT#3 from the BS chamber so that it could have rubber vertical dampers installed, and be characterized.  For future reference, the #'s of the Tip Tilts refers to the serial number of the suspension block piece, which forms the top horizontal bar of the frame. 

ELOG V3.1.3-